OpenStudy (anonymous):

There is a Saturday performance at a local theater. A theater charges $9 for an adult ticket and $4 for a child ticket. All 1020 tickets were sold and $7680 of revenue was brought in. How many adult tickets were sold?

OpenStudy (nurali):

9x+(1020-x)4 = 7680

OpenStudy (redneckoutlaw):

Let x tickets sold were of adults, then (1024-x) were of childrens now, 9x+(1024-x)4 = 7680 above equation is for total money solve it you will get x=720 hence,no of adult tickets were 720 and no. of child tickets were 300

OpenStudy (nurali):

9x+(1020-x)4 = 7680 9x+4080-4x=7680 5x=3600 x=720 hence,no of adult tickets were 720 and no. of child tickets were 300

OpenStudy (redneckoutlaw):

^^

OpenStudy (anonymous):

do i add the tickets together?

OpenStudy (redneckoutlaw):

No, if you did it would say "How many tickets in total were sold" But its just askng how many adult tickets were sold

OpenStudy (anonymous):

ohhh